OpenStudy (anonymous):

An athlete can burn 10 calories per minute while running and 4 calories per minute while walking. If the athlete burns 450 calories in 60 minutes, how long is spent on each activity?

OpenStudy (anonymous):

w = minutes running r = minutes running You can get two equations from the information given: 10r + 4w = 450; // total calories burned = 450 r + w = 60 // total minutes = 60 it follows that: r = 60 - w 10r + 4w = 450 10(60 - w) + 4w = 450 600 - 10w + 4w = 450 600 - 6w = 450 -6w = -150 w = 25 plug into r + w = 60 to get r r + 25 = 60 r = 35 so: w = 25 r = 35